OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin five pi divided by twelve

OpenStudy (briensmarandache):

do you mean sin squared \[\sin ^{2}\left( \frac{ 5\Pi }{ 12 } \right)= .5\left( 1-\cos(2\left( \frac{ 5\Pi }{ 12 } \right) \right)\]

OpenStudy (briensmarandache):

i think that is the way to go about it, sorry it took so long

OpenStudy (briensmarandache):

first time using the equation tab

OpenStudy (briensmarandache):

there is soposed to be another " ) " at the end but you get im sure

OpenStudy (briensmarandache):

supposed

OpenStudy (anonymous):

its not sin^2 its \[\sin \frac{ 5\pi }{ 12 }\] @briensmarandache

OpenStudy (anonymous):

@campbell_st can you help?

OpenStudy (anonymous):

\[\sin\frac{ 5\pi}{ 12} =(+ or -)\sqrt{\frac{ 1+\cos(\frac{ 5\pi }{ 12 } }{ 2}}\]

OpenStudy (anonymous):

doesnt it have to be one half of the 5pi/12 because its the half angle identity

OpenStudy (anonymous):

True

OpenStudy (anonymous):

So 5pi/24

OpenStudy (anonymous):

i think so... now what?

OpenStudy (anonymous):

@swagmaster47 what do i do now?

OpenStudy (anonymous):

Uhm

OpenStudy (anonymous):

\[\sin(5\pi/12) = (+or-)\frac{ \sqrt{1-\cos(5\pi/6)}}{2}\] which equals

OpenStudy (anonymous):

Since sine is positive in quadrant 1 \[\frac{ \sqrt{1+\sqrt(3)/(2)}}{2}\]

OpenStudy (anonymous):

I used 5pi/6 as theta since it is an half angle problem

OpenStudy (anonymous):

And 5pi/6 is in quadrant 1

OpenStudy (anonymous):

my answers can be : \[1/2 \sqrt{2+\sqrt{3}}\] \[1/2 \sqrt{2-\sqrt{3}}\] \[\sqrt{2+\sqrt{3}}\] \[\sqrt{2-\sqrt{3}}\]

OpenStudy (anonymous):

@swagmaster47

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

@SolomonZelman what do you think?

OpenStudy (anonymous):

My interne cut out, here's my process \[\frac{ \sqrt{1+\sqrt(3)/(2)} }{ 2 } = \sqrt{2+\sqrt(3)/(4)} = 1/2 \sqrt{2 + \sqrt(3}\]

OpenStudy (anonymous):

*internet